OpenStudy (anonymous):

algebra help, will give best answer if you answer correctly Which expresses the inequality x ≥ -4 in interval notation? A. (-∞ , -4] B. (-4, ∞ ) C. -4, ∞ ) D. (-∞ , -4)

OpenStudy (anonymous):

Is there supposed to be a bracket on C.? making it [-4, ∞)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

C

OpenStudy (anonymous):

Ok, then the correct answer would be C.

OpenStudy (anonymous):

@segundop thank you for pointing that out. and c looks like the correct answer

OpenStudy (anonymous):

thanks I still need more help 1. Draw on a line graph for the solution set of the inequality x - 4 ≤ 3 2. Which represents the compound inequality in interval notation? x ≥ -3 and x < 1 A. x e (-∞, -3] B. x e (-3, 1) C. x e [-3, 1) D. x e (1, ∞) 3. Draw the solution set of the compound inequality on a line graph x - 2 ≤ 0 or x - 1 ≥ 3

OpenStudy (anonymous):

this is because you want your x value to be greater and or equal to -4, which means that you must include the [ for the -4 and it goes to infinity since there are an infinite amount of numbers that are greater than -4

OpenStudy (anonymous):

thanks segundop, I've rated you as best response. You've been very helpful.

OpenStudy (anonymous):

No prob, ill try helping you out with that second question you posted.

OpenStudy (anonymous):

okay :)

OpenStudy (anonymous):

For this one it is telling you to graph a line graph, first you want to get your x alone on the right side, to do this add both sides by 4, getting you this:\[x \le 7\]

OpenStudy (anonymous):

so it would be 7 going to the left but would the dot be filled in or not filled in?

OpenStudy (anonymous):

yes, the dot would be filled, since \[\le \] means less than or equal to, whenever there is a line underneath the greater sign or less than sign it means that the value can also be equals

OpenStudy (anonymous):

If there is no line underneath the greater than (>) or less than (<) it means that you exclude the value.

OpenStudy (anonymous):

oh okay I get it now

OpenStudy (anonymous):

For example:\[x \ge 7\] In this case you would use the filled in dot to include 7 in the linegraph. \[x >7\] For this case you would not include the value since you are only asked to inclue anything greater than 7 without including 7

OpenStudy (anonymous):

oh okay, what about number 2?

OpenStudy (anonymous):

so with a line it would be filled in, without a dot it wouldn't be filled in

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

okay I understand that now. do you understand #2 and #3?

OpenStudy (anonymous):

yeah, for #2 you have two interval notations, \[x \ge 3 \] and\[x<1\]

OpenStudy (anonymous):

Oops there should be a -3 there: \[x \ge -3\]this means that your x can be great than or equal to -3

OpenStudy (anonymous):

\[x<1\] means that your x value has to be less than 1 (without including 1 because there is no line)

OpenStudy (anonymous):

so would it be B? I know it either has to be B or C

OpenStudy (anonymous):

it would be C, this is because \[x \ge -3 \] which means that the -3 is included in the linegraph

OpenStudy (anonymous):

oh okay that makes sense

OpenStudy (anonymous):

attempt number #3 and post it, ill help me as much as i can